7 Replies Latest reply on Mar 25, 2019 5:06 AM by whiteh

    Unidenitfiable issue on a newly installed 7890B FID

    whiteh

      Hi everyone,

       

      A few weeks ago in our lab we installed a new inlet/column/detector set-up on an existing 7890B GC in our lab. This GC already had a front inlet/detector installed and there was never anything installed on the back components. Therefore we installed a SSL inlet & EPC, capillary column (part # 123-1063), FID detector, 7693 ALS and the 150 sample tray from completely new parts and tied to the existing gas lines.

       

      For the most part, the installation went well and I did the installation myself (I've been to Agilent training courses on GC maintenance so this wasn't a stretch for me). Initially we had some problems with the FID signal spiking a lot, which we resolved once by removing the jet and cleaning it. After this I was able to run about 6 samples which looked good and ran with no issue.

       

      Then we started seeing the signal increase rapidly as soon as the oven temperature was adjusted. For example, if you manually changed the oven set point from the 50C to any number, the second you pressed enter the signal would climb into the thousands and wouldn't come down even if you left it to "bakeout" for some time. Our first assumptions was more detector issues or some kind of electronics issues. We tinkered around as much as we could (see below) and then got in touch with an Agilent technical support rep. Since talking with the Agilent rep we've been able to change it from the signal shooting up quickly to where the signal is still increasing, but it holds off on the sharp increase until about 260C and doesn't reach the thousands.

       

      We're at a stalemate with the instrument and nothing we've done has moved us past the current issue. I've attached an annotated screenshot of the chromatograph and instrument actuals that shows what we see when manually heating the oven (the issue also occurs if running a sample in a proper run). I'm curious if anyone has seen a similar issue before and might know what to do? At this point we have replaced/changed everything except the electronics and for that we're going to need an Agilent visit. We would like to avoid a visit if its some glaringly obvious issue that we're just overlooking. Thanks in advance.


      Everything we've done:

      1. Confirm the issue is likely not associated with the detector

           -With the FID functioning and the flame lit we have used a blanking plug to isolate the detector from the column and repeated the process of manually heating the oven and see no change on the chromatograph

           -We have swapped columns and the issue continues no matter which column is installed

           -We use helium as a carrier gas currently and decided tie in a hydrogen line to see if we maybe had a carrier gas contamination issue, but this didn't help either. Our helium/hydrogen/air cylinders are connected to the other 5 GC instruments in our labs and none of those are experiencing issues.

           -To try and rule out a bad electrical connection we went ahead and pulled apart, then re-plugged in all the connections for the new set up but this didn't help

           -We've had the ALS & sample tray taken off the instrument the whole time we have been troubleshooting, so I don't think the issue could be with either of those.

       

      2. Suspect that the inlet is the cause

           -If we have the inlet at pressure and temperature and we start heating the oven (and consequently see the signal increase) the moment we turn off the inlet pressure the signal begins dropping. If you turn off the pressure and go so far as to remove the septum nut and septum to quickly remove the pressure we see the signal decrease immediately back to the normal baseline

       

      3. Change every inlet consumable possible

           -Setpum, liner, o-ring, gold seal, split vent line, and the split vent cartridge have all been swapped, and none of shown any change in the issue

       

      4. Swap the whole inlet out

           - The last ditch effort was to remove the inlet & its EPC and replace it with another inlet/EPC from an instrument in our lab that we know works. This also did not help.

       

      Any insight would be appreciated before we schedule an Agilent rep to come check it out.